Calculate the volume of a given mass of gas at 760mmhg when its volume was increased by 10dm3 st 710 mmHg of the gas

1
Expert's answer
2022-02-19T06:51:14-0500

Let the volume of the gas be "V1"

So "V2" "=V1+10dm^3"

"P1V1=P2V2"

"760\u00d7V1=710\u00d7(V1+10dm^3)"

"760V1=710(V1+10dm^3)"

"760V1=710V1+7100"

"760V1-710V1=7100"

"50V1=7100"

"\frac{50V1}{50}=\frac{7100}{50}=142"

"V1=142dm^3"